OpenStudy (anonymous):

compute a value for y that satisfies the equation below -3y-12=-4

OpenStudy (anonymous):

Add 12 both the sides first..

OpenStudy (anonymous):

add 12 on both sides. then divide by -3 on both sides.

OpenStudy (anonymous):

-3y-12+12=-4+12 -3y=8 -3y/-3=8/-3 y=-(8/3)

OpenStudy (anonymous):

add 12 to the both sides -3y-12+12=-4+12 -3y=8 divide both sides by -3 y=-8/3

OpenStudy (anonymous):

Not given just a single second to solve the asker on his\her own...

OpenStudy (anonymous):

i get 0.375 is this wrong because i dont know how you get 8/3

OpenStudy (anonymous):

add 12 to bothe sides and you get -3y=8 then you divide 8 by -3 and get...

OpenStudy (anonymous):

is the answer 2 2/3

OpenStudy (anonymous):

y=-2.67

OpenStudy (anonymous):

If you get that answer... Then can you please show your work @naomie_dacelin

OpenStudy (anonymous):

i got it now i divided wrong

OpenStudy (anonymous):

good:D

OpenStudy (anonymous):

\(\large \frac{8}{3}\) can be written as : \(\large 2 \frac{2}{3}\)

OpenStudy (anonymous):

Yeah that right

OpenStudy (anonymous):

but it would be negative

OpenStudy (anonymous):

sorry my mistake

OpenStudy (anonymous):

so what is the real answer